# Scaling of poroelastic coarsening and elastic arrest in crosslinked gels

### v1

@@ V Acknowledgements
Appendix A Onuki-Puri numerical calculations
A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction. In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus. They found that for an uncrosslinked gel (their effective crosslink density, is equivalent to the modulus $G$ in units of $k_{B}T$) coarsening was governed by a 1/3 power law. When varying degrees of crosslinking were included, the coarsening was governed by an effective power law where the exponent depended on the crosslink density (see their Fig. 2 and caption). They also observed the onset of elastic arrest with a time that depended on the value of the modulus. The dynamics of their model only included polymer-solvent friction and the gel had no viscoelastic regime and could not include Darcy transport, poroelastic stress diffusion, and viscoelastic stress buildup, all important elements of our approach and its relevance to the experiments in [9]. In their simulations, the effective network–solvent friction includes an explicit Stokes prefactor ($6\pi$), which enters directly into the arrest criterion. As a result, the arrested length can be enhanced by roughly an order of magnitude relative to the naive elastocapillary estimate $\sigma/G$, even in the absence of an additional large control parameter. By contrast, in our scaling theory the viscosities enter only through ratios (see the Tables below), so such numerical prefactors cancel. The remaining microscopic input enters the arrest length only to a fractional power, rendering the predicted mesoscopic magnification robust and weakly sensitive to microscopic details.
Appendix B Curk and Luijten model
A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\sigma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix. Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects, and does not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach (see Appendix for further discussion). They predict an initial regime of growth followed by elastic arrest at short times, an intermediate viscoelastic coarsening regime with linear-in-time growth, and a late-time crossover back to LSW coarsening with domain size scaling as $t^{1/3}$. The late-time return to LSW scaling is not observed in PDMS gels, where the long-time (hours) state is one of elastic arrest, indicating a distinct rate-limiting physical mechanism.
Appendix C Unified diffusion-based description of coarsening in fluids and gels
@@ C.2 Mobility in binary fluids and gels
where $c_{B}$ is the number density of the transported species. Equivalently, using $\nabla\mu=\nabla p/c_{B}$, the flux may be written as $J\sim-D\nabla c_{B}$. In this case, the mobility is due to the rapid diffusive motion of individual molecules in the surrounding fluid.
@@
The capillary driving force originates in the curvature dependence of the chemical potential at the interface. For a spherical domain of radius $\lambda$, the chemical potential at the interface exceeds that of a flat interface by
$\delta\mu=\frac{\sigma}{c\,\lambda},$ (17)
where $\sigma$ is the interfacial tension and $c$ is the concentration of the transported species in the domain.
@@ C.3 Relation of mechanical and osmotic pressures
By contrast, at sufficiently long times the network responds elastically over length scales comparable to the domain size, and elastic stresses accumulate over these large scales. In this elastic regime, further solvent transport requires deformation of the entire domain, and the resulting pressure gradients are determined by force balance across the domain rather than by local osmotic gradients alone. This buildup of elastic stress leads to elastic arrest of coarsening.
Thus, while elastic stresses are essential for determining the arrest of coarsening at long times, on the coarsening time scale itself the polymer network behaves viscoelastically, and elastic stresses relax locally and do not contribute significantly to the pressure gradients that drive flow.
Capillary stresses enter separately as boundary conditions at interfaces: the Laplace law fixes the pressure jump across the interface between the solvent-rich domain and the gel and thereby sets the magnitude of the pressure difference that must be accommodated by the pressure field in the surrounding material.
In summary, while the mechanical pressure field includes osmotic, elastic, and capillary contributions, gradients of chemical potential may be equivalently expressed in terms of pressure gradients because osmotic stresses provide the fundamental thermodynamic driving force, with transient and local elastic stresses controlling how this driving force is transmitted in the viscoelastic regime of the gel.
@@
$\Pi=\mu\,c_{B}-f(c_{B}).$
To make this relation explicit, we differentiate
$\Pi=\mu\,c-f(c)$
with respect to the state variables at fixed temperature. Using the chain rule, this gives
@@
Substituting this expression into the equation above, the terms proportional to $dc$ cancel, leaving
@@
This shows that gradients in chemical potential are equivalently expressed as gradients in osmotic pressure, which provide the mechanical driving force for transport. Therefore, at fixed temperature,
@@
We recall that $c$ is the concentration of the transported species (the B molecules in the AB small-molecule mixture or the solvent molecules in the solvent-gel system) outside the coarsening solvent-rich (or B molecule rich) domain.
@@
$p(r=\lambda,t)=\delta p\equiv\frac{\sigma}{\lambda}.$ (22)
C.5 Diffusion equation for the pressure field
@@ C.6 Exact solution around a spherical domain
This expression shows explicitly how the interfacial pressure gradient evolves in time as the pressure field spreads into the surrounding medium.
@@ C.9 Coarsening laws
For a binary fluid of small molecules, transport in the continuous phase is due to ordinary molecular diffusion with diffusion constant $D$. As shown above, the pressure field outside a shrinking domain satisfies a diffusion equation and, because molecular diffusion is fast compared to the coarsening dynamics, the pressure profile rapidly reaches its long-time limit. Equivalently, the pressure satisfies Laplace’s equation outside the domain, with the capillary boundary condition $p(r=\lambda)=\delta p\sim\sigma/\lambda$.
@@
$\left.\nabla p\right|_{r=\lambda}\sim\frac{\delta p}{\lambda}\sim\frac{\sigma}{\lambda^{2}}.$
@@
$\dot{\lambda}\sim-\frac{1}{\Delta c}\,\frac{M}{c}\left.\nabla p\right|_{r=\lambda}\sim-\frac{1}{\Delta c}\,D\,\frac{\sigma}{\lambda^{2}}.$ (39)
Integrating this equation yields the standard LSW coarsening law $\lambda(t)\sim(D\sigma\,t/\Delta c)^{1/3}$.
C.9.2 Gel (poroelastic; slow poroelastic diffusion, $\ell_{p}\ll\lambda$).
The interfacial pressure is set by capillarity, $\delta p\sim\sigma/\lambda$, and Darcy flow gives $M\sim k_{0}c_{s}/\eta_{D}$ with $\nabla\mu=(1/c_{s})\nabla p$. Heree, $c_{s}$ is the solvent concentration in the gel where it is transported by coarsening of the solvent-rich domain. For slow poroelastic spreading, the exact diffusion solution yields $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}\sim(\sigma/\lambda)/\ell_{p}$, so that
$\dot{\lambda}\sim-\frac{1}{\Delta c_{s}}\,\frac{k_{0}c_{s}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\frac{c_{s}}{\Delta c_{s}}\,\frac{\sigma}{\lambda\,\ell_{p}(t)},\qquad\ell_{p}(t)=\sqrt{D_{p}t},\ \ D_{p}\sim\frac{k_{0}G}{\eta_{ps}}.$ (40)
@@
C.9.3 Gel (poroelastic; fast poroelastic diffusion, $\ell_{p}\gg\lambda$).
Once poroelastic diffusion has spread over distances comparable to the domain size, the interfacial gradient saturates to $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda\sim\sigma/\lambda^{2}$, and the evolution law becomes
$\dot{\lambda}\sim-\,\frac{k_{0}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\,\frac{\sigma}{\lambda^{2}}.$ (41)
@@
Depending on whether the interfacial pressure gradient has reached its saturated value $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda$ or is still building up in time as $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}(t)$, one recovers either the standard LSW coarsening law or the poroelastic coarsening laws discussed in the main text.

### v2

@@ VI Acknowledgements
Appendix A Curk and Luijten model
A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\gamma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix. Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects, and does not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach (see Appendix for further discussion). They predict an initial regime of growth followed by elastic arrest at short times, an intermediate viscoelastic coarsening regime with linear-in-time growth, and a late-time crossover back to LSW coarsening with domain size scaling as $t^{1/3}$. The late-time return to LSW scaling is not observed in PDMS gels, where the long-time (hours) state is one of elastic arrest, indicating a distinct rate-limiting physical mechanism.
Appendix B Onuki-Puri numerical calculations
A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction. In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus. They found that for an uncrosslinked gel (their effective crosslink density, is equivalent to the modulus $G$ in units of $k_{B}T$) coarsening was governed by a 1/3 power law. When varying degrees of crosslinking were included, the coarsening was governed by an effective power law where the exponent depended on the crosslink density (see their Fig. 2 and caption). They also observed the onset of elastic arrest with a time that depended on the value of the modulus. The dynamics of their model only included polymer-solvent friction and the gel had no viscoelastic regime and could not include Darcy transport, poroelastic stress diffusion, and viscoelastic stress buildup, all important elements of our approach and its relevance to the experiments in [9]. In their simulations, the effective network–solvent friction includes an explicit Stokes prefactor ($6\pi$), which enters directly into the arrest criterion. As a result, the arrested length can be enhanced by roughly an order of magnitude relative to the naive elastocapillary estimate $\gamma/G$, even in the absence of an additional large control parameter. By contrast, in our scaling theory the viscosities enter only through ratios (see the Tables below), so such numerical prefactors cancel. The remaining microscopic input enters the arrest length only to a fractional power, rendering the predicted mesoscopic magnification robust and weakly sensitive to microscopic details.
Appendix C Unified diffusion-based description of coarsening in fluids and viscoelastic gels
@@ C.2 Mobility in binary fluids and gels
where $c_{B}$ is the number density of the transported species. Equivalently, using $\nabla\mu=\nabla p/c_{B}$ which is appropriate for fluids, the flux may be written as $J\sim-D\nabla c_{B}$. In this case, the mobility is due to the rapid diffusive motion of individual molecules in the surrounding fluid.
@@
In both cases the capillary driving force originates in the curvature dependence of the chemical potential at the interface. For a spherical domain of radius $\lambda$, the chemical potential at the interface exceeds that of a flat interface by
$\delta\mu=\frac{\gamma}{c\,\lambda},$ (18)
where $\gamma$ is the interfacial tension and $c$ is the concentration of the transported species in the domain.
@@ C.3 Relation of mechanical and osmotic pressures
By contrast, at sufficiently long times the network responds elastically over length scales comparable to the domain size, and elastic stresses accumulate over these large scales. In this elastic regime, further solvent transport requires deformation of the entire domain, and the resulting pressure gradients are determined by force balance across the domain rather than by local osmotic gradients alone. This buildup of elastic stress leads to elastic arrest of coarsening (see Appendix D. Thus, while elastic stresses are essential for determining the arrest of coarsening, on the coarsening time scale itself the polymer network behaves viscoelastically, and elastic stresses relax locally and do not contribute significantly to the pressure gradients that drive flow.
Capillary stresses enter separately as boundary conditions at interfaces: the Laplace law fixes the pressure jump across the interface between the solvent-rich domain and the gel and thereby sets the magnitude of the pressure difference that is the boundary condition on the pressure field in the surrounding material.
In summary, while the mechanical pressure field in fluids includes osmotic and capillary contributions, gradients of chemical potential may be equivalently expressed in terms of pressure gradients because osmotic stresses provide the fundamental thermodynamic driving force, with transient and local elastic stresses controlling how this driving force is transmitted (via energy dissipation) in the viscoelastic regime of the gel.
@@
$\Pi=\mu\,c-f(c).$
To make this relation explicit, we differentiate $\Pi$ at fixed temperature. Using the chain rule,
@@
Substituting, the terms proportional to $dc$ cancel, leaving
@@
Therefore, when the free energy depends only on concentration,
@@
Limitation to fluid phases. The proportionality above relies on the assumption that
$f=f(c)$
is a function of concentration alone. In an elastic solid or gel, however, the free-energy density depends on additional independent variables such as the strain $\varepsilon$:
$f=f(c,\varepsilon).$
In that case,
$df=\left(\frac{\partial f}{\partial c}\right)_{\varepsilon}dc+\left(\frac{\partial f}{\partial\varepsilon}\right)_{c}d\varepsilon,$
and the chemical potential is defined at fixed strain,
$\mu=\left(\frac{\partial f}{\partial c}\right)_{\varepsilon}.$
Repeating the derivation gives
$d\Pi=c\,d\mu-\left(\frac{\partial f}{\partial\varepsilon}\right)_{c}d\varepsilon.$
Thus, in an elastic medium where strain varies spatially, gradients in chemical potential are not simply proportional to gradients in pressure. The additional mechanical term couples composition and deformation, so transport cannot in general be written purely in terms of $\nabla\Pi$. The relation $\nabla\Pi=c\,\nabla\mu$ holds only for systems whose thermodynamics is fully described by concentration alone (e.g. two-component fluids), but not for elastic systems in which strain is an independent degree of freedom. In the relation of the osmotic pressure and the chemical potential, $c$ denotes the concentration of the transported species (the B molecules in the AB small-molecule mixture or the solvent molecules in the solvent–gel system) outside the coarsening solvent-rich (or B-rich) domain.
@@
$p(r=\lambda,t)=\delta p\equiv\frac{\gamma}{\lambda}.$ (23)
C.5 Diffusion equation for the pressure field outside the coarsening domain
@@ C.6 Exact solution around a spherical domain
where $\delta p\sim\gamma/\lambda$. This expression shows explicitly how the interfacial pressure gradient evolves in time as the pressure field spreads into the surrounding medium.
@@ C.9 Coarsening laws
For a binary fluid of small molecules, transport in the continuous phase is due to ordinary molecular diffusion with diffusion constant $D$. As shown above, the pressure field outside a shrinking domain satisfies a diffusion equation and, because molecular diffusion is fast compared to the coarsening dynamics, the pressure profile rapidly reaches its long-time limit. Equivalently, the pressure satisfies Laplace’s equation outside the domain, with the capillary boundary condition $p(r=\lambda)=\delta p\sim\gamma/\lambda$.
@@
$\left.\nabla p\right|_{r=\lambda}\sim\frac{\delta p}{\lambda}\sim\frac{\gamma}{\lambda^{2}}.$
@@
$\dot{\lambda}\sim-\frac{1}{\Delta c}\,\frac{M}{c}\left.\nabla p\right|_{r=\lambda}\sim-\frac{1}{\Delta c}\,D\,\frac{\gamma}{\lambda^{2}}.$ (40)
Integrating this equation yields the standard LSW coarsening law $\lambda(t)\sim(D\,\gamma\,t/\Delta c)^{1/3}$.
C.9.2 Gel (poroelastic; slow poroelastic diffusion)
The interfacial pressure is set by capillarity, $\delta p\sim\gamma/\lambda$, and Darcy flow gives $M\sim k_{0}c_{s}/\eta_{D}$ with $\nabla\mu=(1/c_{s})\nabla p$. Heree, $c_{s}$ is the solvent concentration in the gel where it is transported by coarsening of the solvent-rich domain. For slow poroelastic spreading, the exact diffusion solution yields $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}\sim(\gamma/\lambda)/\ell_{p}$, so that
$\dot{\lambda}\sim-\frac{1}{\Delta c_{s}}\,\frac{k_{0}c_{s}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\frac{c_{s}}{\Delta c_{s}}\,\frac{\gamma}{\lambda\,\ell_{p}(t)},\qquad\ell_{p}(t)=\sqrt{D_{p}t},\ \ D_{p}\sim\frac{k_{0}G}{\eta_{ps}}.$ (41)
@@ C.9.2 Gel (poroelastic; slow poroelastic diffusion)
C.9.3 Gel (poroelastic; fast poroelastic diffusion)
Once poroelastic diffusion has spread over distances comparable to the domain size, the interfacial gradient saturates to $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda\sim\gamma/\lambda^{2}$ (see Eq. 35), and the evolution law becomes
$\dot{\lambda}\sim-\,\frac{k_{0}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\,\frac{\gamma}{\lambda^{2}}.$ (42)
@@ C.9.3 Gel (poroelastic; fast poroelastic diffusion)
Depending on whether the interfacial pressure gradient has reached its saturated value $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda$ or is still building up in time as $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}(t)$, with $\ell_{p}=\sqrt{D_{p}\,t}$, one recovers either the standard LSW coarsening law $t^{1/3}$ or the slow poroelastic coarsening law $t^{1/4}$ discussed in the main text.
Appendix D Elastic stress boundary condition for a spherical liquid domain in an elastic, infinite and incompressible gel
We derive the elastic stress at the boundary of a spherical solvent-rich domain of radius $\lambda$ embedded in an infinite, incompressible linear elastic gel of shear modulus $G$, when the interface is displaced radially by a small amount $\delta\lambda$. The interface is characterized by an isotropic surface tension $\gamma$.
Throughout this section we emphasize an important physical point: in a gel the solvent inside the domain and the solvent in the surrounding porous network are the same continuous phase. Therefore, in mechanical equilibrium the hydrostatic solvent pressure must be spatially uniform. Only pressure gradients drive Darcy flow.
D.1 Displacement, strain and stress
The spherical solvent-rich domain occupies $r<\lambda$ and the elastic gel occupies $r>\lambda$. By spherical symmetry, the displacement in the gel is purely radial,
$\mathbf{u}(\mathbf{x})=u(r)\,\hat{\mathbf{r}},\qquad r>\lambda.$ (43)
For small strains, the only nonzero strain components are
$\varepsilon_{rr}=\frac{du}{dr},\qquad\varepsilon_{\theta\theta}=\varepsilon_{\phi\phi}=\frac{u}{r}.$ (44)
Incompressibility of the gel (polymer network + solvent mixture) requires
$\mathrm{tr}\,\varepsilon=\varepsilon_{rr}+2\varepsilon_{\theta\theta}=\frac{du}{dr}+2\frac{u}{r}=0.$ (45)
Thus
$\frac{d}{dr}\left(r^{2}u\right)=0\quad\Rightarrow\quad u(r)=\frac{A}{r^{2}},\qquad r>\lambda.$ (46)
For an incompressible, isotropic, linear elastic gel, the polymer stress tensor is written
$\bm{\sigma}=-p(r)\,\mathbf{I}+2G\,\bm{\varepsilon}.$ (47)
Here $p(r)$ is the isotropic pressure (often called the pore pressure in poroelasticity). Mathematically, it appears as a Lagrange multiplier enforcing incompressibility. Physically, it may be identified (up to an arbitrary additive constant) with the hydrostatic pressure of the solvent. Only its gradient has physical consequences for solvent flow through Darcy’s law. The pressure that enters the polymer stress tensor must equal the solvent pressure in mechanical equilibrium for an incompressible system.
D.2 Boundary conditions and displacement field
We impose a prescribed radial displacement at the interface:
$u(\lambda)=\delta\lambda.$ (48)
and determine $\delta\lambda$ below in the arrested state. For a free-standing gel, we use stress-free conditions at infinity:
$\bm{\sigma}(r\to\infty)\to\mathbf{0}.$ (49)
[…]
$\delta\lambda=\frac{A}{\lambda^{2}}\quad\Rightarrow\quad A=\delta\lambda\,\lambda^{2}.$ (50)
Hence
$u(r)=\delta\lambda\left(\frac{\lambda}{r}\right)^{2}.$ (51)
[…]
ε_θθ(r) =ε_ϕϕ(r) =δλλ2r3. One immediately verifies incompressibility:
$\varepsilon_{rr}+2\varepsilon_{\theta\theta}=0.$
[…]
σ_θθ(r) = -p(r) +2G δλλ2r3. Mechanical equilibrium requires that the divergence of the stress tensor vanish and in spherical coordinates this means:
$\frac{d\sigma_{rr}}{dr}+\frac{2}{r}\left(\sigma_{rr}-\sigma_{\theta\theta}\right)=0.$ (52)
Substitution gives
$\frac{dp}{dr}=0,$ (53)
so the pressure gradient is zero and we set
$p(r)=p_{\infty},$ (54)
a spatial constant. Since the deviatoric stresses vanish at infinity, the stress-free boundary condition at infinity implies $p_{\infty}=0$. This choice fixes a constant pressure term: only pressure differences matter physically.
D.3 Interfacial normal stress balance and continuity of solvent pressure
The normal stress exerted by the gel at the interface is
$\sigma_{rr}^{\text{gel}}(\lambda)=-4G\,\frac{\delta\lambda}{\lambda},$ (55)
while inside the solvent-rich domain the stress (negative of pressure within the solvent) is purely hydrostatic,
$\bm{\sigma}^{(s)}=-p_{s}\,\mathbf{I}.$
An elastic, permanently crosslinked network can coexist with solvent flow (poroelasticity). In that case, quasi-static momentum balance for the polymer skeleton includes the interphase drag:
$\nabla\!\cdot\bm{\sigma}^{(p)}+\bm{f}_{\mathrm{drag}}=\mathbf{0},\qquad\bm{f}_{\mathrm{drag}}=\frac{\eta}{k}\,(\bm{v}_{s}-\bm{v}_{p}).$ (56)
Therefore, “mechanical equilibrium” does not by itself imply $\nabla\!\cdot\bm{\sigma}^{(p)}=\mathbf{0}$ unless $\bm{f}_{\mathrm{drag}}=\mathbf{0}$.
However, in the arrested elastic regime considered here, the permanently crosslinked polymer network is static in the bulk and experiences no distributed (bulk) body forces. Surface tension provides a typical capillary stress scale $\sim\gamma/\lambda$ (with $\lambda$ a typical radius of curvature / domain size), but it enters the problem only through interfacial traction (normal-stress) boundary conditions. In an elastic solid, these interfacial tractions can be balanced locally by the network’s deviatoric (non-isotropic) elastic stresses, rather than by a spatially varying pressure field of the solvent in the pores of the gel. Under this arrested/static assumption, the polymer bulk equilibrium (away from the boundaries) reduces to the zero net-force condition: $\nabla\!\cdot\bm{\sigma}^{(p)}=\mathbf{0}$.
[…]
The Young–Laplace traction condition reads
$\sigma_{rr}^{\text{gel}}(\lambda)-\sigma_{rr}^{(\ell)}(\lambda)=\frac{2\gamma}{\lambda}.$ (57)
Using $\sigma_{rr}^{(\ell)}=-p_{\ell}=-p_{\infty}=0$, this reduces to
$\sigma_{rr}^{\text{gel}}(\lambda)=\frac{2\gamma}{\lambda}.$ (58)
Substituting the elastic stress gives
$4G\,\frac{\delta\lambda}{\lambda}=\frac{2\gamma}{\lambda},$ (59)
which determines the displacement $\delta\lambda$
$\delta\lambda\sim\frac{\gamma}{G}$ (60)
(up to a numerical prefactor depending on sign convention).
To summarize: In elastic equilibrium, the isotropic solvent pressure is spatially uniform so no Darcy flow occurs. The Laplace normal stress is balanced entirely by deviatoric (non-isotropic) elastic stress. This is fundamentally different from the viscoelastic regime of the gel, where the Laplace pressure appears as a jump in solvent pressure which then decreases as a power law within the gel. That solvent pressure gradient drives Darcy flow of the solvent from the solvent-rich domain into the gel, i.e., coarsening. In contrast, in the elastic regime, the solvent (and polymer) pressure remains uniform and the force that balances the capillary (surface tension) force is elastic and deviatoric. The lack of a solvent pressure gradient means that Darcy flow of the solvent, and hence coarsening is arrested.